OpenStudy (anonymous):

The world record for the 100 m sprint is held by Jamaica's Usain Bolt at 9.58 (a) At this speed how long does it take bolt to run 10 m ? how do I solve it if the answer is 9.69?

OpenStudy (unklerhaukus):

\[s=\frac dt\]\[t=\frac ds\] \[s_{\text{Bolt}}=\frac{100[\text m]}{9.58[\text s]}\] \[t_{\text{Bolt}}=\frac d{s_{\text{Bolt}}}=\frac{10[\text m]}{{100[\text m]}/{9.58[\text s]}}=\qquad\dots[\text s]\]